# Analysis of Shuffling Beyond Pure Local Differential Privacy

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
Shuffling has emerged as one of the most powerful tools for improving the utility–privacy trade-off of differentially private (DP) distributed data analysis (Liew et al., 2022; Luo et al., 2025). In the single-message model, a population of users first applies a local randomizer to their data and then submits the resulting messages to an anonymizing shuffler. A growing body of work has demonstrated that, even when the local randomizer satisfies only relatively weak privacy guarantees, shuffling can significantly amplify privacy at the level of the aggregated mechanism, leading to substantially improved accuracy compared with pure local DP (Feldman et al., 2022; Balle et al., 2019; Erlingsson et al., 2019).
From both a theoretical and a practical perspective, two complementary ingredients are crucial. On the one hand, we need general analytic results that explain how and to what extent shuffling amplifies privacy for a given local mechanism. On the other hand, practitioners ultimately require numerically accurate finite-$n$ estimates of the privacy parameters $(\varepsilon,\delta)$ for concrete protocols. In the central-DP setting, the latter need has led to a shift from coarse composition theorems to tight numerical accountants (for example, for DP-SGD), which in turn has enabled substantially stronger utility–privacy trade-offs in practice (Gopi et al., 2021). In contrast, shuffle DP currently lacks, both from a theoretical and a practical perspective, tools that offer a comparable level of accuracy and flexibility.
@@ 1. Introduction
Most existing theoretical and numerical analyses of shuffling focus on local randomizers that satisfy pure local DP (i.e., $(\varepsilon_{0},0)$-local DP) and treat $\varepsilon_{0}$ as the main knob controlling privacy amplification (see, e.g., (Balle et al., 2019; Erlingsson et al., 2019)). This $\varepsilon_{0}$-centric viewpoint has two important limitations.
First, the parameter $\varepsilon_{0}$ summarizes the local privacy guarantee, but it is a rather crude descriptor from the viewpoint of shuffle DP: it largely ignores the structural properties of the local mechanism that actually govern privacy amplification. In particular, existing general bounds typically do not distinguish between, say, a Laplace mechanism with a given noise scale, a $k$-randomized response ($k$-$R$) mechanism with a given $k$, or other natural families of local mechanisms. As a consequence, generic upper bounds expressed solely in terms of $\varepsilon_{0}$ can be quite loose. Indeed, several works have shown that by exploiting the specific structure of a given mechanism one can obtain significantly tighter bounds on the effect of shuffling (e.g., for $k$-$R$ one can show that privacy amplification becomes stronger as $k$ increases, even when $\varepsilon_{0}$ is kept fixed (Feldman et al., 2022; Biswas et al., 2024)). Yet, we currently lack a general understanding of what intrinsic properties of a local mechanism govern its “shuffle efficiency.”
Second, the restriction to pure local DP mechanisms excludes a wide range of natural and practically relevant local randomizers. Pure local DP is a convenient sufficient condition for privacy amplification by shuffling, but it is far from necessary. In practice, one often encounters approximate-DP local mechanisms or mechanisms that do not satisfy any DP condition. Surprisingly, even for the Gaussian mechanism, arguably one of the most prominent mechanisms in central DP, it is not known how to precisely characterize privacy amplification by shuffling. Balle et al. (Balle et al., 2019) already pointed out that the Gaussian case is technically challenging, and subsequent work by Koskela et al. (Koskela et al., 2022) observed that even numerical accounting for Gaussian shuffling is nontrivial. To date, existing results for shuffling of Gaussian mechanisms are essentially limited to lower bounds (Liew and Takahashi, 2022; Chua et al., 2024). While there are attempts to generalize shuffling guarantees to approximate local DP (i.e., $(\varepsilon_{0},\delta_{0}>0)$-local DP) in a mechanism-agnostic fashion (Feldman et al., 2022; Cheu et al., 2019), such general bounds tend to be overly pessimistic; they often show little visible amplification for moderate values of $n$, because they are driven by worst-case mechanisms that do not reflect the structure of the specific mechanism.
@@
A central tool in the analysis of shuffle DP is the privacy blanket introduced by Balle et al. (Balle et al., 2019). Given a local randomizer $\mathcal{R}$, the privacy blanket is a data-independent distribution $\mathcal{R}_{\mathrm{BG}}$ defined using $\mathcal{R}$. Balle et al. showed that $\delta$ of the shuffled mechanism can be upper-bounded by a divergence computed under the blanket distribution; we refer to this quantity as the blanket divergence. This blanket-divergence bound is currently the best-known general upper bound on privacy amplification by shuffling (Su et al., 2025).
However, the blanket divergence itself is still poorly understood, largely due to its analytical and numerical intractability. In particular, Balle et al. (Balle et al., 2019) mainly derive generic upper bounds via Hoeffding- and Bennett-type concentration inequalities based on the local DP parameter $\varepsilon_{0}$. Consequently, these bounds can be quite loose and fail to capture how the blanket divergence depends on the structure of the underlying local randomizer, as well as how tight the blanket-based upper bounds are compared with known lower bounds.
@@
Another line of work seeks to go beyond the $\varepsilon_{0}$-centric view by exploiting additional structure of the local randomizer, typically in the form of a small number of scalar parameters. Feldman et al. (Feldman et al., 2023) introduced a stronger clone paradigm and showed that, for mechanisms such as $k$-$R$, one can parameterize the effect of shuffling using $\varepsilon_{0}$ together with two additional scalar quantities, yielding substantially tighter bounds than $\varepsilon_{0}$-only analyses, although even in the $k$-$R$ case exact optimality is not established. Building on this, Wang et al. (Wang et al., 2024, 2025) refine the stronger clone-based analysis, using two scalar parameters, and provide sufficient conditions on the local randomizer under which their bounds become optimal. A subtle point is that the proof of Lemma 4.5 in (Wang et al., 2024) implicitly assumes the existence of a nontrivial post-processing map (see the errata of (Feldman et al., 2023)). Outside specific mechanism classes such as $k$-$R$ where no such map is needed (see Theorem 7.1 of (Feldman et al., 2023)), the results should therefore be interpreted with some care.
Several works have proposed numerical methods for evaluating shuffle DP guarantees. Koskela et al. (Koskela et al., 2022) use the clone paradigm together with FFT to approximate the privacy loss of shuffled mechanisms. Their approach, however, is restricted to pure local DP mechanisms and inherits the looseness of the clone paradigm, as shown by Su et al. (Su et al., 2025). Su et al. (Su et al., 2025) propose a numerical analysis of the blanket divergence based on FFT, similar in spirit to ours. However, they still assume pure local DP and do not provide rigorous control of the relative error; in addition, the running time scales as $O(n^{2})$ in the worst case when targeting a constant additive error. Biswas et al. (Biswas et al., 2024) give an exact combinatorial method for computing the privacy guarantees of shuffled $k$-$R$, but their analysis focuses on fixed neighboring database pairs rather than general $(\varepsilon,\delta)$-DP guarantees.
@@ 2.1. Differential Privacy (DP)
Definition 2.2 (Differential privacy via hockey-stick divergence (Dwork and Roth, 2013; Barthe and Olmedo, 2013)).
@@
Sometimes, we say that $\mathcal{M}$ satisfies $(\varepsilon,\delta)$-DP for $x_{1:n}\simeq x^{\prime}_{1:n}$ if $\mathcal{D}_{e^{\varepsilon}}\bigl(\mathcal{M}(x_{1:n})\,\big\|\,\mathcal{M}(x^{\prime}_{1:n})\bigr)\leq\delta$ (Sommer et al., 2019).
@@
Following Balle et al. (Balle et al., 2019), we work in the single-message randomize-then-shuffle model.That is, we assume black-box access to an ideal functionality shuffler [^2] In contrast to the multi-message shuffling model, which typically requires access to a stronger ideal functionality $\mathcal{S}$ capable of splitting or decorrelating multiple messages from the same user (for example by hiding timing and grouping information), the single-message shuffling model only relies on a comparatively simple “anonymization” ideal functionality that permutes user identities. This makes the single-message model significantly easier to implement in practice, and therefore in this work we restrict attention to the single-message shuffling setting..
@@
Lemma 2.5 (Lemma 5.3 of Balle et al. (Balle et al., 2019)).
@@
  • If $\mathcal{R}_{\mathrm{ref}}=\mathcal{R}_{\mathrm{BG}}$ is the blanket distribution, then $l_{\varepsilon}(Y)$ is the privacy amplification random variable underlying the blanket-divergence upper bound of Balle et al. (Balle et al., 2019).
  • If $\mathcal{R}_{\mathrm{ref}}=\mathcal{R}_{x}$ for some $x\in\mathcal{X}$, then $l_{\varepsilon}(Y)$ is the privacy amplification random variable appearing in the lower bound of Su et al. (Su et al., 2025).
2.2. Local Randomizers Beyond Pure Local DP
@@
Accordingly, we formulate our results in a mechanism-agnostic way and do not commit to a specific local randomizer. Nevertheless, it is useful to keep in mind a concrete running example satisfying Assumption 2.8 without pure local DP and truncation. In particular, we will often refer to the generalized Gaussian mechanism, which forms a flexible family of location mechanisms (parameterized by the shape $\beta$) and simultaneously encompasses the Laplace mechanism ($\beta=1$) and the Gaussian mechanism ($\beta=2$). The latter case is especially noteworthy, as the Gaussian mechanism has been explicitly highlighted in prior work (e.g., Balle et al. (Balle et al., 2019) and Koskela et al. (Koskela et al., 2022)) as technically challenging to analyze in the shuffle model.
@@
3. Analysis of Shuffling via the Central Limit Theorem (CLT)
@@
3.1. Analysis of Blanket Divergence
@@
Lemma 3.1 (Transformations for the blanket divergence).
@@
(1) $\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})$ $=\frac{1}{\gamma}\,\mathbb{E}\Bigl[Z_{1}\,\Pr\bigl[S_{n-1}>-Z_{1}\bigr]\Bigr]$ $=\frac{1}{\gamma}\,\mathbb{E}\!\left[Z_{1}\,\bigl\{1-F_{n-1}\bigl(t_{n-1}-\eta_{n-1}(Z_{1})\bigr)\bigr\}\right].$
(2) $\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})$ $=\frac{1}{n\gamma}\int_{0}^{\infty}\Pr[S_{n}>u]\,du$ $=\frac{\sigma_{\varepsilon}}{\sqrt{n}\gamma}\,\int_{t_{n}}^{\infty}\bigl(1-F_{n}(x)\bigr)\,dx.$
@@
Lemma 3.2 (Asymptotic expansion of the blanket divergence).
@@
(3) $\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon_{n}},n,\mathcal{R}_{\mathrm{ref}},\gamma}\;=\;\varphi\bigl(\chi\bigl(\mathrm{e}^{\varepsilon_{n}}-1\bigr)\sqrt{n}\bigr)\,\left(\frac{1}{\chi^{3}}\,\frac{1}{\bigl(\mathrm{e}^{\varepsilon_{n}}-1\bigr)^{2}n^{3/2}}\right)\bigl(1+o(1)\bigr)\;=\;\varphi\bigl(\chi\varepsilon_{n}\sqrt{n}\bigr)\,\left(\frac{1}{\chi^{3}}\,\frac{1}{\varepsilon_{n}^{2}n^{3/2}}\right)\bigl(1+o(1)\bigr).$
@@ 3.3. Shuffle Index
We defer the proof of Theorem 3.4 to Appendix A.4. In particular, the theorem applies to the case of $k$-$R$ ($k\geq 3$), and hence to mechanisms that are built on top of $k$-$R$ or share the same blanket structure, such as local hashing (Wang et al., 2017) and (for sufficiently small caps) PrivUnit (Asi et al., 2022). Thus, in many practically relevant shuffle protocols, the blanket-divergence analysis is asymptotically optimal in the sense of Theorem 3.4. For more complex families such as generalized Gaussian mechanisms, the upper and lower shuffle indices do not coincide in general, but across a wide range of parameters we empirically observe that their ratio $\underline{\chi}_{\mathrm{lo}}/\underline{\chi}_{\mathrm{up}}$ remains above $0.7$. The detailed calculations of the shuffle indices and their numerical values are deferred to Appendix B.
4. Analysis of Blanket Divergence via the Fast Fourier Transform (FFT)

### v2

@@ Abstract
1 Introduction
Shuffling has emerged as one of the most powerful tools for improving the utility–privacy trade-off of differentially private (DP) distributed data analysis Liew et al. (2022); Luo et al. (2025). In the single-message model, a population of users first applies a local randomizer to their data and then submits the resulting messages to an anonymizing shuffler. A growing body of work has demonstrated that, even when the local randomizer satisfies only relatively weak privacy guarantees, shuffling can significantly amplify privacy at the level of the aggregated mechanism, leading to substantially improved accuracy compared with pure local DP Feldman et al. (2022); Balle et al. (2019); Erlingsson et al. (2019).
From both a theoretical and a practical perspective, two complementary ingredients are crucial. On the one hand, we need general analytic results that explain how and to what extent shuffling amplifies privacy for a given local mechanism. On the other hand, practitioners ultimately require numerically accurate finite-$n$ estimates of the privacy parameters $(\varepsilon,\delta)$ for concrete protocols. In the central-DP setting, the latter need has led to a shift from coarse composition theorems to tight numerical accountants (for example, for DP-SGD), which in turn has enabled substantially stronger utility–privacy trade-offs in practice Gopi et al. (2021). In contrast, shuffle DP currently lacks, both from a theoretical and a practical perspective, tools that offer a comparable level of accuracy and flexibility.
@@ 1 Introduction
Most existing theoretical and numerical analyses of shuffling focus on local randomizers that satisfy pure local DP (i.e., $(\varepsilon_{0},0)$-local DP) and treat $\varepsilon_{0}$ as the main knob controlling privacy amplification (see, e.g., Balle et al. (2019); Erlingsson et al. (2019)). This $\varepsilon_{0}$-centric viewpoint has two important limitations.
First, the parameter $\varepsilon_{0}$ summarizes the local privacy guarantee, but it is a rather crude descriptor from the viewpoint of shuffle DP: it largely ignores the structural properties of the local mechanism that actually govern privacy amplification. In particular, existing general bounds typically do not distinguish between, say, a Laplace mechanism with a given noise scale, a $k$-randomized response ($k$-$R$) mechanism with a given $k$, or other natural families of local mechanisms. As a consequence, generic upper bounds expressed solely in terms of $\varepsilon_{0}$ can be quite loose. Indeed, several works have shown that by exploiting the specific structure of a given mechanism one can obtain significantly tighter bounds on the effect of shuffling (e.g., for $k$-$R$ one can show that privacy amplification becomes stronger as $k$ increases, even when $\varepsilon_{0}$ is kept fixed Feldman et al. (2022); Biswas et al. (2024)). Yet, we currently lack a general understanding of what intrinsic properties of a local mechanism govern its “shuffle efficiency.”
Second, the restriction to pure local DP mechanisms excludes a wide range of natural and practically relevant local randomizers. Pure local DP is a convenient sufficient condition for privacy amplification by shuffling, but it is far from necessary. In practice, one often encounters approximate-DP local mechanisms or mechanisms that do not satisfy any DP condition. Surprisingly, even for the Gaussian mechanism, arguably one of the most prominent mechanisms in central DP, it is not known how to precisely characterize privacy amplification by shuffling. Balle et al. Balle et al. (2019) already pointed out that the Gaussian case is technically challenging, and subsequent work by Koskela et al. Koskela et al. (2022) observed that even numerical accounting for Gaussian shuffling is nontrivial. To date, existing results for shuffling of Gaussian mechanisms are essentially limited to lower bounds Liew and Takahashi (2022); Chua et al. (2024). While there are attempts to generalize shuffling guarantees to approximate local DP (i.e., $(\varepsilon_{0},\delta_{0}>0)$-local DP) in a mechanism-agnostic fashion Feldman et al. (2022); Cheu et al. (2019), such general bounds tend to be overly pessimistic; they often show little visible amplification for moderate values of $n$, because they are driven by worst-case mechanisms that do not reflect the structure of the specific mechanism.
@@
A central tool in the analysis of shuffle DP is the privacy blanket introduced by Balle et al. Balle et al. (2019). Given a local randomizer $\mathcal{R}$, the privacy blanket is a data-independent distribution $\mathcal{R}_{\mathrm{BG}}$ defined using $\mathcal{R}$. Balle et al. showed that $\delta$ of the shuffled mechanism can be upper-bounded by a divergence computed under the blanket distribution; we refer to this quantity as the blanket divergence. This blanket-divergence bound is currently the best-known general upper bound on privacy amplification by shuffling Su et al. (2025).
However, the blanket divergence itself is still poorly understood, largely due to its analytical and numerical intractability. In particular, Balle et al. Balle et al. (2019) mainly derive generic upper bounds via Hoeffding- and Bennett-type concentration inequalities based on the local DP parameter $\varepsilon_{0}$. Consequently, these bounds can be quite loose and fail to capture how the blanket divergence depends on the structure of the underlying local randomizer, as well as how tight the blanket-based upper bounds are compared with known lower bounds.
@@
Another line of work seeks to go beyond the $\varepsilon_{0}$-centric view by exploiting additional structure of the local randomizer, typically in the form of a small number of scalar parameters. Feldman et al. Feldman et al. (2023) introduced a stronger clone paradigm and showed that, for mechanisms such as $k$-$R$, one can parameterize the effect of shuffling using $\varepsilon_{0}$ together with two additional scalar quantities, yielding substantially tighter bounds than $\varepsilon_{0}$-only analyses, although even in the $k$-$R$ case exact optimality is not established. Building on this, Wang et al. Wang et al. (2024, 2025) refine the stronger clone-based analysis, using two scalar parameters, and provide sufficient conditions on the local randomizer under which their bounds become optimal. A subtle point is that the proof of Lemma 4.5 in Wang et al. (2024) implicitly assumes the existence of a nontrivial post-processing map (see the errata of Feldman et al. (2023)). Outside specific mechanism classes such as $k$-$R$ where no such map is needed (see Theorem 7.1 of Feldman et al. (2023)), the results should therefore be interpreted with some care.
Several works have proposed numerical methods for evaluating shuffle DP guarantees. Koskela et al. Koskela et al. (2022) use the clone paradigm together with FFT to approximate the privacy loss of shuffled mechanisms. Their approach, however, is restricted to pure local DP mechanisms and inherits the looseness of the clone paradigm, as shown by Su et al. Su et al. (2025). Su et al. Su et al. (2025) propose a numerical analysis of the blanket divergence based on FFT, similar in spirit to ours. However, they still assume pure local DP and do not provide rigorous control of the relative error; in addition, the running time scales as $O(n^{2})$ in the worst case when targeting a constant additive error. Biswas et al. Biswas et al. (2024) give an exact combinatorial method for computing the privacy guarantees of shuffled $k$-$R$, but their analysis focuses on fixed neighboring database pairs rather than general $(\varepsilon,\delta)$-DP guarantees.
@@ 2.1 Differential Privacy (DP)
Definition 2.2 (Differential privacy via hockey-stick divergence Dwork and Roth (2013); Barthe and Olmedo (2013)).
@@
Sometimes, we say that $\mathcal{M}$ satisfies $(\varepsilon,\delta)$-DP for $x_{1:n}\simeq x^{\prime}_{1:n}$ if $\mathcal{D}_{e^{\varepsilon}}\bigl(\mathcal{M}(x_{1:n})\,\big\|\,\mathcal{M}(x^{\prime}_{1:n})\bigr)\leq\delta$ Sommer et al. (2019).
@@
Following Balle et al. Balle et al. (2019), we work in the single-message randomize-then-shuffle model.That is, we assume black-box access to an ideal functionality shuffler [^2] In contrast to the multi-message shuffling model, which typically requires access to a stronger ideal functionality $\mathcal{S}$ capable of splitting or decorrelating multiple messages from the same user (for example by hiding timing and grouping information), the single-message shuffling model only relies on a comparatively simple “anonymization” ideal functionality that permutes user identities. This makes the single-message model significantly easier to implement in practice, and therefore in this work we restrict attention to the single-message shuffling setting..
@@
Remark 2.1 (Lemma 5.3 of Balle et al. Balle et al. (2019)).
@@
  • If $\mathcal{R}_{\mathrm{ref}}=\mathcal{R}_{\mathrm{BG}}$ is the blanket distribution, then $l_{\varepsilon}(Y)$ is the privacy amplification random variable underlying the blanket-divergence upper bound of Balle et al. Balle et al. (2019).
  • If $\mathcal{R}_{\mathrm{ref}}=\mathcal{R}_{x}$ for some $x\in\mathcal{X}$, then $l_{\varepsilon}(Y)$ is the privacy amplification random variable appearing in the lower bound of Su et al. Su et al. (2025).
2.2 Local Randomizers Beyond Pure Local DP
@@
Accordingly, we formulate our results in a mechanism-agnostic way and do not commit to a specific local randomizer. Nevertheless, it is useful to keep in mind a concrete running example satisfying Assumption 2.3 without pure local DP and truncation. In particular, we will often refer to the generalized Gaussian mechanism, which forms a flexible family of location mechanisms (parameterized by the shape $\beta$) and simultaneously encompasses the Laplace mechanism ($\beta=1$) and the Gaussian mechanism ($\beta=2$). The latter case is especially noteworthy, as the Gaussian mechanism has been explicitly highlighted in prior work (e.g., Balle et al. Balle et al. (2019) and Koskela et al. Koskela et al. (2022)) as technically challenging to analyze in the shuffle model.
@@
3 Analysis of Shuffling via the Central Limit Theorem (CLT)
@@
3.1 Analysis of Blanket Divergence
@@
Remark 3.1 (Transformations for the blanket divergence).
@@
$\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})$ $=\frac{1}{\gamma}\,\mathbb{E}\Bigl[Z_{1}\,\Pr\bigl[S_{n-1}>-Z_{1}\bigr]\Bigr]$ $=\frac{1}{\gamma}\,\mathbb{E}\!\left[Z_{1}\,\bigl\{1-F_{n-1}\bigl(t_{n-1}-\eta_{n-1}(Z_{1})\bigr)\bigr\}\right].$ (1)
$\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})$ $=\frac{1}{n\gamma}\int_{0}^{\infty}\Pr[S_{n}>u]\,du$ $=\frac{\sigma_{\varepsilon}}{\sqrt{n}\gamma}\,\int_{t_{n}}^{\infty}\bigl(1-F_{n}(x)\bigr)\,dx.$ (2)
@@
Remark 3.2 (Asymptotic expansion of the blanket divergence).
@@
$\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon_{n}},n,\mathcal{R}_{\mathrm{ref}},\gamma}\;=\;\varphi\bigl(\chi\bigl(\mathrm{e}^{\varepsilon_{n}}-1\bigr)\sqrt{n}\bigr)\,\left(\frac{1}{\chi^{3}}\,\frac{1}{\bigl(\mathrm{e}^{\varepsilon_{n}}-1\bigr)^{2}n^{3/2}}\right)\bigl(1+o(1)\bigr)\;=\;\varphi\bigl(\chi\varepsilon_{n}\sqrt{n}\bigr)\,\left(\frac{1}{\chi^{3}}\,\frac{1}{\varepsilon_{n}^{2}n^{3/2}}\right)\bigl(1+o(1)\bigr).$ (3)
@@ 3.3 Shuffle Index
We defer the proof of Theorem 3.4 to Appendix A.4. In particular, the theorem applies to the case of $k$-$R$ ($k\geq 3$), and hence to mechanisms that are built on top of $k$-$R$ or share the same blanket structure, such as local hashing Wang et al. (2017) and (for sufficiently small caps) PrivUnit Asi et al. (2022). Thus, in many practically relevant shuffle protocols, the blanket-divergence analysis is asymptotically optimal in the sense of Theorem 3.4. For more complex families such as generalized Gaussian mechanisms, the upper and lower shuffle indices do not coincide in general, but across a wide range of parameters we empirically observe that their ratio $\underline{\chi}_{\mathrm{lo}}/\underline{\chi}_{\mathrm{up}}$ remains above $0.7$. The detailed calculations of the shuffle indices and their numerical values are deferred to Appendix B.
4 Analysis of Blanket Divergence via the Fast Fourier Transform (FFT)